find the length of the major axis containing the foci of the ellipse whose equation is 10x^2 + 20y^2 = 200 in simplest radical form.
Answer by lwsshak3(11628)

find the length of the major axis containing the foci of the ellipse whose equation is 10x^2 + 20y^2 = 200 in simplest radical form.
**
Standard form of an equation for and ellipse with horizontal major axis:
(x-h)^2/a^2+(y-k)^2/b^2=1, a>b, (h,k) being the (x,y) coordinates of the center.
..
10x^2 + 20y^2 = 200
divide by 200
x^2/20+y^2/10=1
center: (0,0)
a^2=20
a=√20
length of horizontal major axis=2a=2√20